5x + 7y = 19 find three solutions

Answer by JayJay_T(6)

Try to find a x-value at which the 19-5x part can easily be divided by 7:
5x + 7y = 19
=> 7y = 19 - 5x
So suppose x = -6
=> 7y = 19 - (5.-6)
=> 7y = 19 - (-30)
=> 7y = 19 - -30
=> 7y = 19 + 30
=> 7y = 49
=> y = 49/7
=> y = 7
Solution: (-6,7)
Other solutions via the same principle: (-12,11) and (8,-3)